# TalkingHeadBench: A Multi-Modal Benchmark & Analysis of Talking-Head DeepFake Detection

### v2

@@ 4.1 Evaluation Setup
SOTA Detectors To address these questions, we benchmark seven publicly available SOTA deepfake detectors: CADDM [dong2023implicit], LipFD [liu2024lips], DeepFake-Adapter [shao2025deepfake], TALL [xu2023tall], AltFreezing [wang2023altfreezing], MM-Det [song2025learningmultimodalforgeryrepresentation], and HiFi-Net [hifi_net_xiaoguo]. Similar to the generators, we aim to diversify the modalities and frameworks of our detectors, as shown in Tab. 3. These seven detectors meet our goal of being the latest SOTA detectors with available code, and together, they capture varying modalities.
@@
For these protocols, we select 1600 real videos from FaceForensics++ [rossler2019faceforensics++] and CelebV-HQ [zhu2022celebv], unseen during generation, to match fake video numbers per protocol, and perform face comparison to prevent identity leakage between train and test across real and fake video sets.
Evaluation Metrics Detector performance under these protocols is measured with three metrics: AUC, Brier Score, TPR@FPR=1% (T1), and TPR@FPR=0.1% (T0.1), following IJB-C face benchmark [maze2018iarpa]. Stricter thresholds on FPR, e.g., T0.1, are more useful when these detectors are applied at scale and only a small amount of bad detections are tolerable. We will prioritize T1 for most analysis in this paper, following its use in the IJB-C face benchmark, but later we will analyze the detector performance across various thresholds in FPR. We classify the performance of detectors as:
@@ 4.2 Impact of Generator and Identity Shifts
Protocol-Dependent Performance Shifts: Overall, detectors exhibit performance shifts across protocols. In particular, average $AUC$ and T1 scores tend to decline from P1 to P2, indicating that generator shift (P2) poses greater challenges than identity shift (P1). This trend continues in P3, where combined generator and identity shifts further reduce performance as illustrated in Fig. 2a. For instance, TALL’s [xu2023tall] average T1 score decreases from 0.99 in P1 to 0.93 in P2 and 0.87 in P3. Despite these increasingly difficult conditions, TALL and DeepFake-Adapter [shao2025deepfake] maintain strong performance across all three protocols, with DeepFake-Adapter’s lower score on P1 largely due to poor performance on AniPortrait Audio [wei2024aniportrait](T1 = 0.58; see Tab. 5(a)). This underscores the need to assess detectors not just by averages, but also by generator-level sensitivity.
Generator Difficulty and Cross-Protocol Generalization: Fig. 2b shows the variation of detector performance across generators. EMOPortraits [drobyshev2024emoportraits] stands out as the most challenging case, with average T1 of only 0.44, far below all others. By contrast, the remaining generators cluster within a narrower range of 0.60-0.69, indicating more comparable levels of detectability. This sharp gap emphasizes that while detectors maintain moderate robustness across most generators, EMOPortraits remains a persistent failure case that reveals their limitations under stricter thresholds.
@@ 4.3 Detector Robustness and SOTA Evaluation
Overall Robustness Profile: While detailed per protocol analysis is valuable, evaluating robustness across thresholds provides a comprehensive view of a detector’s generalization capabilities. Fig. 3 summarizes performance across all generators, averaged over protocols, for T1 and T0.1 thresholds. High and balanced coverage across all axes indicates consistent generalization across identity, generator, and joint shifts—crucial for handling unseen data from future SOTA generators and varying identities. At T1 (Fig. 3a), DeepFake-Adapter [shao2025deepfake] and TALL [xu2023tall] exhibit strong performance across most generators, with DeepFake-Adapter excelling on LivePortrait [guo2024liveportrait], Hallo [xu2024hallo], and AniPortraitVideo [wei2024aniportrait]. However, when the threshold is tightened to T0.1 (Fig. 3b), DeepFake-Adapter’s performance drops sharply—particularly for AniPortraitAudio [wei2024aniportrait] and EMOPortraits [drobyshev2024emoportraits]—indicating reduced robustness under stricter constraints. In contrast, TALL retains high scores across most generators at T0.1, suggesting better threshold-level resilience and stronger generalization under a low false positive rate. We also evaluate detector performance generalizing to unseen generators such as commercial generator MAGI-1 [magi1] and academic generator Hallo3 [cui2024hallo3]. We observe similar trend whereas the best-performing models, TALL and DeepFake-Adapter, maintain high detection scores. We include the full result in the Supplementary.
@@ 4.4 Generalization to Emerging Generators
To simulate a real-world scenario, we evaluated detectors against deepfakes from emerging academic generator Hallo3 [cui2024hallo3] and commercial generator MAGI-1 [magi1], which were not seen during training. The results, provided in supplementary (Section 4), show that top-performing models like TALL and AltFreezing maintained high accuracy, suggesting that training on the diverse and curated set in TalkingHeadBench enables detectors to handle novel generator architectures. This indicates the detectors have learned a fundamental representation of talking-head artifacts beyond overfitting to specific generator fingerprints.
@@ 4.5 Error Analysis and Explainability
The quantitative results highlight areas requiring deeper investigation, for instance, the poor performance of most detectors against EMOPortraits [drobyshev2024emoportraits], particularly in P3 where T1 scores drop significantly. Similarly, CADDM [dong2023implicit] underperforms consistently on TalkingHeadBench, indicating a need to examine its limitations more closely.
We leverage explainability techniques like Grad-CAM [Selvaraju_2019] to visualize the regions and features that detectors focus their predictions on. We demonstrate an example for TALL [xu2023tall] since it achieves the overall best performance on our dataset. Other detectors are in the supplementary.
Failure Case Analysis: On the EMOPortraits dataset in P3-TALL [xu2023tall]’s worst performing scenario (see Tab. 5(c)), failure cases reveal that the model misclassifies both real and fake videos due to reliance on background cues. As shown in Fig. 5, real and fake videos are getting misclassified due to the model’s attention being disproportionately focused on the background, despite meaningful facial cues suggesting the correct label. This indicates a weakness in TALL’s generalization: it is overly influenced by background artifacts, likely due to training bias, and fails to localize attention to the face under combined identity and generator shifts.
@@
Success Case Analysis: In contrast to its failures, TALL demonstrates effective generalization in certain EMOPortraits [drobyshev2024emoportraits] videos under P3, as shown in Fig. 4. The model accurately classifies a deepfake by focusing on the neck region, which is known to exhibit artifacts in EMOPortraits. Notably, this success occurs despite such artifacts being uncommon in other datasets used during training. This suggests TALL [xu2023tall] has transferred knowledge from broader manipulation cues across datasets. The focused activation around the lower face and neck reflects meaningful model behavior and highlights its potential for cross-domain generalization.
@@
EMOPortraits [drobyshev2024emoportraits] 188 72 38.30%
Hallo2 [cui2024hallo2] 102 9 8.82%
LivePortrait [guo2024liveportrait] 168 5 2.98%
@@ 4.6 Guidance for Future Research using TalkingHeadBench
Limitations of SOTA Detectors. Our analysis shows that while models like TALL [xu2023tall] and DeepFake-Adapter [shao2025deepfake] perform strongly across most generators, there remains room for improvement in achieving consistent excellence across domains. Specifically, TALL achieves exceptional T1 scores on 3 of 6 generators and no worse than fair on the others, while DeepFake-Adapter reaches exceptional on 2 and stays within reasonable to fair elsewhere. Notably, both models are jointly exceptional only on AniPortraitVideo [wei2024aniportrait] and Hallo [xu2024hallo], suggesting detectors still struggle to generalize uniformly across diverse generation methods.
Open Challenges. While our analysis offers insights into robustness, several aspects of TalkingHeadBench remain difficult. These include (i) handling generators like EMOPortraits [drobyshev2024emoportraits] and Hallo2 [cui2024hallo2], where SOTA detectors reach only 70% and 88% T1 in P3; (ii) reducing degradation in P3 where both identity and generator differ between splits, as all detectors except DF-Adapter degrade significantly; (iii) overcoming reliance on background artifacts—as shown in Sec. 4.5—which causes models to ignore relevant facial features. These challenges highlight the need for detectors with stronger domain adaptation, semantic attention, and modality-aware training to handle real-world deepfakes more reliably.
@@
(a) Performance: The best model, TALL [xu2023tall], reaches excellent T1 ($99\%$) on only 3/6 generators in P3, while reaching $\sim$90% for Hallo2 and LivePortrait and $\sim$80% for EMOPortraits. Future research should strive for $99\%$ in both T1 and T0.1 across all 6 detectors and protocols.
@@ 5 Conclusions, Broad Impacts, and Limitations
We introduce TalkingHeadBench, a comprehensive benchmark utilizing contemporary talking-head generators to stress-test deepfake detectors under realistic distribution shifts. Our evaluations reveal critical gaps: even state-of-the-art detectors struggle with generalization across changes in identity and generator, particularly with challenging fakes like EMOPortraits [drobyshev2024emoportraits], often due to over-reliance on non-facial artifacts. TalkingHeadBench will broadly impact the community by providing a standardized, challenging benchmark that reflects modern deepfake realism and complexity. We will provide open access to the dataset, evaluation protocols, and baseline code.
@@
The dataset only contains a finite number of videos and identities derived from FFHQ [karras2019style] and CelebV-HQ [zhu2022celebv], which might not represent the diversity of real-world deepfakes that can be exhibited. While we focus on creating high-quality forgeries, we acknowledge that our dataset can inherit the demographic imbalances from the public source datasets. We provide a detailed demographic audit in the Supplementary for transparency, and identify notable disparities across racial and age groups, while gender bias is minimal. In the future, we will set up leaderboards that allow performance comparison among newly developed generators and detectors. Additionally, we will conduct regular fairness audits, applying both group-based and individual fairness metrics to ensure robustness in our future leaderboard as new generators emerge. This benchmark serves as a crucial resource for developing trustworthy deepfake detection systems, essential for mitigating the increasing societal risks posed by manipulated media.
@@ B.1 Source Datasets for Generation and Real Videos
  • CelebV-HQ [zhu2022celebv]: This large-scale video facial attributes dataset was used as the source for dynamic driving signals. It contains a variety of identities, ethnicities, expressions, and poses. Both video segments (for video-driven generators) and extracted audio tracks (for audio-driven generators, converted to .wav format) were utilized. The dataset can be accessed through the following link: https://github.com/CelebV-HQ/CelebV-HQ. The copyright information for this dataset is not explicitly mentioned.
@@
  • FaceForensics++ (FF++) [rossler2019faceforensics++]: Original, unmanipulated video sequences were included in our set of real videos. We directly downloaded the dataset following the download script they provided, and finalized the number of our YouTube real videos from 1000 to 704 due to the unavailability of some of the videos on YouTube. We then renamed the downloaded audio to match the naming of the YouTube real videos. The renamed audio files are in our Hugging Face datasets (./audio/ff++). The original FF++ dataset can be accessed through the following link: https://github.com/ondyari/FaceForensics. The copyright information can be found at https://github.com/ondyari/FaceForensics/blob/master/LICENSE.
  • CelebV-HQ [zhu2022celebv]: To diversify the identity in our real videos, we incorporated a distinct set of real videos from CelebV-HQ that were not used for driving signal extraction.
@@ B.3 DeepFake Detectors
DeepFake-Adapter [shao2025deepfake]
@@ C.1 Dataset Generation, Splits, and Curation Details
Approximately 500-600 videos were initially generated for each of the six open-source academic models. This was achieved by scripting random pairings of source images from FFHQ [karras_style-based_2019] with driving signals (audio or video) from distinct splits of CelebV-HQ [zhu2022celebv].
@@ E.1 MAGI-1 Commercial Generator Results
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 CADDM [dong2023implicit] 0.88 0.29 0.29 0.95 0.39 0.39 0.91 0.53 0.53 0.96 0.62 0.62 0.97 0.62 0.62 0.97 0.64 0.64 HiFi-Net [hifi_net_xiaoguo] 0.66 0.02 0.00 0.89 0.39 0.23 0.85 0.04 0.02 0.73 0.02 0.00 0.86 0.06 0.04 0.87 0.09 0.02 TALL [xu2023tall] 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 DF-Adapter [shao2025deepfake] 1.00 1.00 1.00 1.00 0.99 0.99 1.00 0.99 0.99 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00
@@ E.2 Hallo3 Academic Generator Results
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 CADDM [dong2023implicit] 0.67 0.24 0.24 0.70 0.40 0.40 0.93 0.76 0.76 0.82 0.64 0.64 0.97 0.91 0.91 0.93 0.69 0.69 HiFi-Net [hifi_net_xiaoguo] 0.84 0.19 0.18 0.88 0.30 0.27 0.91 0.24 0.20 0.65 0.24 0.12 1.00 0.99 0.99 0.87 0.02 0.00 TALL [xu2023tall] 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 DF-Adapter [shao2025deepfake] 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 0.99 0.99

### v3

@@ 4.1 Evaluation Setup
SOTA Detectors To address these questions, we benchmark seven publicly available SOTA deepfake detectors: CADDM [10], LipFD [20], DeepFake-Adapter [32], TALL [40], AltFreezing [36], MM-Det [35], and HiFi-Net [14]. Similar to the generators, we aim to diversify the modalities and frameworks of our detectors, as shown in Tab. 3. These seven detectors meet our goal of being the latest SOTA detectors with available code, and together, they capture varying modalities.
@@
For these protocols, we select 1600 real videos from FaceForensics++ [29] and CelebV-HQ [44], unseen during generation, to match fake video numbers per protocol, and perform face comparison to prevent identity leakage between train and test across real and fake video sets.
Evaluation Metrics Detector performance under these protocols is measured with three metrics: AUC, Brier Score, TPR@FPR=1% (T1), and TPR@FPR=0.1% (T0.1), following IJB-C face benchmark [21]. Stricter thresholds on FPR, e.g., T0.1, are more useful when these detectors are applied at scale and only a small amount of bad detections are tolerable. We will prioritize T1 for most analysis in this paper, following its use in the IJB-C face benchmark, but later we will analyze the detector performance across various thresholds in FPR. We classify the performance of detectors as:
@@ 4.2 Impact of Generator and Identity Shifts
Protocol-Dependent Performance Shifts: Overall, detectors exhibit performance shifts across protocols. In particular, average $AUC$ and T1 scores tend to decline from P1 to P2, indicating that generator shift (P2) poses greater challenges than identity shift (P1). This trend continues in P3, where combined generator and identity shifts further reduce performance as illustrated in Fig. 2a. For instance, TALL’s [40] average T1 score decreases from 0.99 in P1 to 0.93 in P2 and 0.87 in P3. Despite these increasingly difficult conditions, TALL and DeepFake-Adapter [32] maintain strong performance across all three protocols, with DeepFake-Adapter’s lower score on P1 largely due to poor performance on AniPortrait Audio [37](T1 = 0.58; see Tab. 5(a)). This underscores the need to assess detectors not just by averages, but also by generator-level sensitivity.
Generator Difficulty and Cross-Protocol Generalization: Fig. 2b shows the variation of detector performance across generators. EMOPortraits [11] stands out as the most challenging case, with average T1 of only 0.44, far below all others. By contrast, the remaining generators cluster within a narrower range of 0.60-0.69, indicating more comparable levels of detectability. This sharp gap emphasizes that while detectors maintain moderate robustness across most generators, EMOPortraits remains a persistent failure case that reveals their limitations under stricter thresholds.
@@ 4.3 Detector Robustness and SOTA Evaluation
Overall Robustness Profile: While detailed per protocol analysis is valuable, evaluating robustness across thresholds provides a comprehensive view of a detector’s generalization capabilities. Fig. 3 summarizes performance across all generators, averaged over protocols, for T1 and T0.1 thresholds. High and balanced coverage across all axes indicates consistent generalization across identity, generator, and joint shifts—crucial for handling unseen data from future SOTA generators and varying identities. At T1 (Fig. 3a), DeepFake-Adapter [32] and TALL [40] exhibit strong performance across most generators, with DeepFake-Adapter excelling on LivePortrait [12], Hallo [39], and AniPortraitVideo [37]. However, when the threshold is tightened to T0.1 (Fig. 3b), DeepFake-Adapter’s performance drops sharply—particularly for AniPortraitAudio [37] and EMOPortraits [11]—indicating reduced robustness under stricter constraints. In contrast, TALL retains high scores across most generators at T0.1, suggesting better threshold-level resilience and stronger generalization under a low false positive rate. We also evaluate detector performance generalizing to unseen generators such as commercial generator MAGI-1 [30] and academic generator Hallo3 [8]. We observe similar trend whereas the best-performing models, TALL and DeepFake-Adapter, maintain high detection scores. We include the full result in the Supplementary.
@@ 4.4 Generalization to Emerging Generators
To simulate a real-world scenario, we evaluated detectors against deepfakes from emerging academic generator Hallo3 [8] and commercial generator MAGI-1 [30], which were not seen during training. The results, provided in supplementary (Section 4), show that top-performing models like TALL and AltFreezing maintained high accuracy, suggesting that training on the diverse and curated set in TalkingHeadBench enables detectors to handle novel generator architectures. This indicates the detectors have learned a fundamental representation of talking-head artifacts beyond overfitting to specific generator fingerprints.
@@ 4.5 Error Analysis and Explainability
The quantitative results highlight areas requiring deeper investigation, for instance, the poor performance of most detectors against EMOPortraits [11], particularly in P3 where T1 scores drop significantly. Similarly, CADDM [10] underperforms consistently on TalkingHeadBench, indicating a need to examine its limitations more closely.
We leverage explainability techniques like Grad-CAM [31] to visualize the regions and features that detectors focus their predictions on. We demonstrate an example for TALL [40] since it achieves the overall best performance on our dataset. Other detectors are in the supplementary.
Failure Case Analysis: On the EMOPortraits dataset in P3-TALL [40]’s worst performing scenario (see Tab. 5(c)), failure cases reveal that the model misclassifies both real and fake videos due to reliance on background cues. As shown in Fig. 5, real and fake videos are getting misclassified due to the model’s attention being disproportionately focused on the background, despite meaningful facial cues suggesting the correct label. This indicates a weakness in TALL’s generalization: it is overly influenced by background artifacts, likely due to training bias, and fails to localize attention to the face under combined identity and generator shifts.
@@
Success Case Analysis: In contrast to its failures, TALL demonstrates effective generalization in certain EMOPortraits [11] videos under P3, as shown in Fig. 4. The model accurately classifies a deepfake by focusing on the neck region, which is known to exhibit artifacts in EMOPortraits. Notably, this success occurs despite such artifacts being uncommon in other datasets used during training. This suggests TALL [40] has transferred knowledge from broader manipulation cues across datasets. The focused activation around the lower face and neck reflects meaningful model behavior and highlights its potential for cross-domain generalization.
@@
EMOPortraits [11] 188 72 38.30%
Hallo2 [7] 102 9 8.82%
LivePortrait [12] 168 5 2.98%
@@ 4.6 Guidance for Future Research using TalkingHeadBench
Limitations of SOTA Detectors. Our analysis shows that while models like TALL [40] and DeepFake-Adapter [32] perform strongly across most generators, there remains room for improvement in achieving consistent excellence across domains. Specifically, TALL achieves exceptional T1 scores on 3 of 6 generators and no worse than fair on the others, while DeepFake-Adapter reaches exceptional on 2 and stays within reasonable to fair elsewhere. Notably, both models are jointly exceptional only on AniPortraitVideo [37] and Hallo [39], suggesting detectors still struggle to generalize uniformly across diverse generation methods.
Open Challenges. While our analysis offers insights into robustness, several aspects of TalkingHeadBench remain difficult. These include (i) handling generators like EMOPortraits [11] and Hallo2 [7], where SOTA detectors reach only 70% and 88% T1 in P3; (ii) reducing degradation in P3 where both identity and generator differ between splits, as all detectors except DF-Adapter degrade significantly; (iii) overcoming reliance on background artifacts—as shown in Sec. 4.5—which causes models to ignore relevant facial features. These challenges highlight the need for detectors with stronger domain adaptation, semantic attention, and modality-aware training to handle real-world deepfakes more reliably.
@@
(a) Performance: The best model, TALL [40], reaches excellent T1 ($99\%$) on only 3/6 generators in P3, while reaching $\sim$90% for Hallo2 and LivePortrait and $\sim$80% for EMOPortraits. Future research should strive for $99\%$ in both T1 and T0.1 across all 6 detectors and protocols.
@@ 5 Conclusions, Broad Impacts, and Limitations
We introduce TalkingHeadBench, a comprehensive benchmark utilizing contemporary talking-head generators to stress-test deepfake detectors under realistic distribution shifts. Our evaluations reveal critical gaps: even state-of-the-art detectors struggle with generalization across changes in identity and generator, particularly with challenging fakes like EMOPortraits [11], often due to over-reliance on non-facial artifacts. TalkingHeadBench will broadly impact the community by providing a standardized, challenging benchmark that reflects modern deepfake realism and complexity. We will provide open access to the dataset, evaluation protocols, and baseline code.
@@
The dataset only contains a finite number of videos and identities derived from FFHQ [18] and CelebV-HQ [44], which might not represent the diversity of real-world deepfakes that can be exhibited. While we focus on creating high-quality forgeries, we acknowledge that our dataset can inherit the demographic imbalances from the public source datasets. We provide a detailed demographic audit in the Supplementary for transparency, and identify notable disparities across racial and age groups, while gender bias is minimal. In the future, we will set up leaderboards that allow performance comparison among newly developed generators and detectors. Additionally, we will conduct regular fairness audits, applying both group-based and individual fairness metrics to ensure robustness in our future leaderboard as new generators emerge. This benchmark serves as a crucial resource for developing trustworthy deepfake detection systems, essential for mitigating the increasing societal risks posed by manipulated media.
@@ B.1 Source Datasets for Generation and Real Videos
  • CelebV-HQ [44]: This large-scale video facial attributes dataset was used as the source for dynamic driving signals. It contains a variety of identities, ethnicities, expressions, and poses. Both video segments (for video-driven generators) and extracted audio tracks (for audio-driven generators, converted to .wav format) were utilized. The dataset can be accessed through the following link: https://github.com/CelebV-HQ/CelebV-HQ. The copyright information for this dataset is not explicitly mentioned.
@@
  • FaceForensics++ (FF++) [29]: Original, unmanipulated video sequences were included in our set of real videos. We directly downloaded the dataset following the download script they provided, and finalized the number of our YouTube real videos from 1000 to 704 due to the unavailability of some of the videos on YouTube. We then renamed the downloaded audio to match the naming of the YouTube real videos. The renamed audio files are in our Hugging Face datasets (./audio/ff++). The original FF++ dataset can be accessed through the following link: https://github.com/ondyari/FaceForensics. The copyright information can be found at https://github.com/ondyari/FaceForensics/blob/master/LICENSE.
  • CelebV-HQ [44]: To diversify the identity in our real videos, we incorporated a distinct set of real videos from CelebV-HQ that were not used for driving signal extraction.
@@ B.3 DeepFake Detectors
DeepFake-Adapter [32]
@@ C.1 Dataset Generation, Splits, and Curation Details
Approximately 500-600 videos were initially generated for each of the six open-source academic models. This was achieved by scripting random pairings of source images from FFHQ [karras_style-based_2019] with driving signals (audio or video) from distinct splits of CelebV-HQ [44].
@@ E.1 MAGI-1 Commercial Generator Results
AniAudio [37] AniVideo [37] Hallo [39] Hallo2 [7] EMOPortraits [11] LivePortrait [12] Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 CADDM [10] 0.88 0.29 0.29 0.95 0.39 0.39 0.91 0.53 0.53 0.96 0.62 0.62 0.97 0.62 0.62 0.97 0.64 0.64 HiFi-Net [14] 0.66 0.02 0.00 0.89 0.39 0.23 0.85 0.04 0.02 0.73 0.02 0.00 0.86 0.06 0.04 0.87 0.09 0.02 TALL [40] 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 DF-Adapter [32] 1.00 1.00 1.00 1.00 0.99 0.99 1.00 0.99 0.99 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00
@@ E.2 Hallo3 Academic Generator Results
AniAudio [37] AniVideo [37] Hallo [39] Hallo2 [7] EMOPortraits [11] LivePortrait [12] Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 CADDM [10] 0.67 0.24 0.24 0.70 0.40 0.40 0.93 0.76 0.76 0.82 0.64 0.64 0.97 0.91 0.91 0.93 0.69 0.69 HiFi-Net [14] 0.84 0.19 0.18 0.88 0.30 0.27 0.91 0.24 0.20 0.65 0.24 0.12 1.00 0.99 0.99 0.87 0.02 0.00 TALL [40] 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 DF-Adapter [32] 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 0.99 0.99